The following frequency table represents the weight of a random sample of 50 students where the largest value is 112 and the smallest value is 44.
a) Find the range and class width.
b) Complete the first 2 rows of the following table.
Class Limit | Class Boundaries | Midpoint | Frequency | Cumulative Frequency | Relative
Frequency
5
6
1
Expert's answer
2020-12-24T14:24:31-0500

a) Find the range and class width.


Range = Largest Value - Smallest Value = 112 - 44 = 88


If the number of classes is 9 then

class width = Range/number of classes = 88/9 = 9.8

but we need to round up, so we will use 10.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 Limit & Boundaries & Midpoint & Frequency & Cum Fr & RelFr \\ \hline
 44-53 & 43.5-53.5 & 48.5 & 5 & 5 & 0.1 \\
 \hdashline
 54-63 & 53.5-63.5 & 58.5 & 6 & 11 & 0.12 \\
\end{array}"